OpenStudy (anonymous):

A radioactive substance decays exponentially. A scientist begins with 130 milligrams of a radioactive substance. After 24 hours, 65 mg of the substance remains. How many milligrams will remain after 45 hours?

OpenStudy (anonymous):

\[N(t) = N * .5^{t/h} \] where n(t) is amount of substance left, N = initial amount, t = current time, and h = half-life time. h = 24 hours t= 45 hours, N = 130mg therefore, \[N(t) = 130 * .5 ^{45/24} = 35.44mg\]

OpenStudy (anonymous):

you are given the half life is 24 hours. so the formula is \[A = 130(\frac{1}{2})^{\frac{t}{24}}\] replace t by 45 daniel has it!

OpenStudy (anonymous):

thank you!:)